OpenStudy (studygurl14):

@mathmale @ganeshie8 @ParthKohli @Nnesha @mathstudent55 @freckles @uri \(\huge\int_{0}^{k} 4kx-5k\ dx\)

OpenStudy (freckles):

k is a constant you can bring it outside the integral

OpenStudy (freckles):

\[k \int\limits_0^k (4x-5) dx=?\]

OpenStudy (freckles):

\[k \int\limits x^n dx=k \frac{x^{n+1}}{n+1}+C , n \neq -1 \\ \]

OpenStudy (studygurl14):

Oh! Thank you!